OpenStudy (anonymous):

A bracelet is marked down 20%. The original price is $37.50. How much is the discount? $7.50 $30.00 $45.00 $1.88

Directrix (directrix):

The discount is 20% of the original price of $37.50.

Directrix (directrix):

% means hundredths. 20% means 20/100 = 1/5 = .2 Discount = .2 * 37.50 = ?